theyluvmark:

A boat traveled 240 miles each way downstream and back. The trip downstream took 10 hours. The trip back took 40 hours. What is the speed of the boat in still water? What is the speed of the current?

theyluvmark:

how

Macho14:

first, divide the miles traveled by the time downstream \[240/10\]

Macho14:

it equals 24

Macho14:

then do the same thing for the time downstream \[240/40\]

Macho14:

this equals 6

Macho14:

add the two values

Macho14:

and divide by 2

Macho14:

30/2 = 15

Macho14:

the boat speed is 15 mph

theyluvmark:

thank you

surjithayer:

let the speed of boat=x m/hr speed of stream=y m/hr time taken downstream=10 hrs time taken upstrestream= 40 hrs resultant speed downstream=(x+y) m/hr resultant speed upstream=(x-y) m/r then (x+y)(10)=240 x+y=24 ...(1) (x-y)(40)=240 x-y=240/40=6 ...(2) add (1) and (2) 2x=30 x=30/2=15 x+y=24 15+y=24 y=24-15=9 speed of boat=15 m/hr speed of stream=9 m/hr

theyluvmark:

@surjithayer wrote:
let the speed of boat=x m/hr speed of stream=y m/hr time taken downstream=10 hrs time taken upstrestream= 40 hrs resultant speed downstream=(x+y) m/hr resultant speed upstream=(x-y) m/r then (x+y)(10)=240 x+y=24 ...(1) (x-y)(40)=240 x-y=240/40=6 ...(2) add (1) and (2) 2x=30 x=30/2=15 x+y=24 15+y=24 y=24-15=9 speed of boat=15 m/hr speed of stream=9 m/hr
thank you as well

Macho14:

sorry, someone called me, I wasn't able to give you the speed of the stream

surjithayer:

yw